OpenStudy (anonymous):

i dont know how to solve this question plz help me: Solve 5^x=4 . What are the exact and approximate solutions?

OpenStudy (lopus):

you must know logarithms:\[\ln (a^{x})= x \ln (a)\]

OpenStudy (lopus):

then apply in both sides ln ln(5^x)=ln(4)